# Language-specific Calibration for Pruning Multilingual Language Models

### v1

@@ Abstract
Recent advances in large language model (LLM) pruning have shown state-of-the-art compression results in post-training and retraining-free settings while maintaining high predictive performance. However, such research mainly considers calibrating pruning using English text, despite the multilingual nature of modern LLMs and their frequent uses in non-English languages. In this paper, we set out to explore effective strategies for calibrating the pruning of multilingual language models. We present the first comprehensive empirical study, comparing different calibration languages for pruning multilingual models across diverse tasks, models, and state-of-the-art pruning techniques. Our results present practical suggestions, for example, calibrating in the target language can efficiently yield lower perplexity, but does not necessarily benefit downstream tasks. Our further analysis experiments unveil that calibration in the target language mainly contributes to preserving language-specific features related to fluency and coherence, but might not contribute to capturing language-agnostic features such as language understanding and reasoning. Last, we provide practical recommendations for future practitioners.
Language-specific Calibration for Pruning Multilingual Language Models
Simon Kurz†∗ Jian-Jia Chen†∗ Lucie Flek⋄∗ Zhixue Zhao‡ †Technical University Dortmund, ⋄University of Bonn, ‡University of Sheffield, ∗Lamarr Institute for Machine Learning and Artificial Intelligence {simon.kurz, jian-jia.chen}@tu-dortmund.de flek@bit.uni-bonn.de zhixue.zhao@sheffield.ac.uk
@@ 1 Introduction
State-of-the-art language models often rely on over-parameterization with millions or billions of parameters, resulting in significant memory and computational demands (Zhang et al., 2017; Allen-Zhu et al., 2019; Xu and Du, 2023). Pruning is a model compression technique that removes unimportant weights to reduce memory footprint and inference computation (Gholami et al., 2022; Hoefler et al., 2021; Kuzmin et al., 2023). Recent pruning methods for language models, such as SparseGPT (Frantar and Alistarh, 2023) and Wanda (Sun et al., 2024), demonstrated state-of-the-art performance in a post-training and retraining-free setting (Zhu et al., 2023). The pruning process involves passing a small number of examples, i.e. calibration data, to the model to determine the importance of weights for subsequent pruning (Kuzmin et al., 2022; Frantar and Alistarh, 2023; Kuzmin et al., 2023).
Notably, existing studies typically use calibration data in English and evaluate the post-pruning performance in English. On the other hand, most state-of-the-art LLMs are multilingual and frequently employed for tasks in non-English languages (Touvron et al., 2023; Achiam et al., 2023; Jiang et al., 2023). It remains unclear how to calibrate pruning to optimize the post-pruning performance of multilingual LLMs on tasks in non-English languages. In other words, the current literature provides little insight into efficient calibration strategies for multilingual LLMs targeting specific non-English languages. For example, if we aim to prune a multilingual LLM and use the pruned LLM for tasks in German, should we calibrate the compression process with text in English, German, or both?
This research, as the first, presents an extensive study investigating the impact of calibration data language on pruning LLMs for language-specific tasks. We recognize the significant value of our study, particularly in resource-constrained scenarios and for languages with limited resources. The main takeaways are the following:
Calibrating on the target language consistently yields the lowest perplexity, but will not guarantee optimal performance on downstream tasks.
Pruning re-orders the strength language of the multilingual model. That is, the model sacrifices performance in some strength languages for others after the pruning.
No single pruning technique consistently outperforms others across different models and tasks. In general, SparseGPT is recommended for pruning Llama-3 8B, while both Wanda and SparseGPT exhibit mixed performance with Aya-23 8B.
Pruning substantially impacts the storage and retrieval of knowledge in a multilingual model across different languages.
@@ 2.1 Multilingual Language Models
Most current state-of-the-art LMs, such as Llama-3 (Meta AI, 2024) and Phi 3(Abdin et al., 2024), are trained on multilingual data, enabling them to understand and generate text in multiple languages (Huang et al., 2023; Holmström et al., 2023; Xu et al., 2024; Meta AI, 2024). Although these multilingual LMs follow the general training paradigm of training monolingual LMs, they are often found to behave differently or show unique features from monolingual models (Xu et al., 2024). For example, Deng et al. (2024) reveals that multilingual LMs are prone to generate unsafe outputs on malicious instructions in non-English languages, i.e. multilingual jailbreak. Wang et al. (2023) also identify that all LLMs tested produce significantly more unsafe responses for non-English queries than English ones. Furthermore, previous work on model explanation finds that multilingual models are different from counterpart monolingual models in terms of their internal processes, that is, different importance distributions on the same input (Jørgensen et al., 2022; Zhao and Aletras, 2024). A general conclusion drawn from these studies is that findings from monolingual settings, particularly those in English, are unlikely to hold in multilingual settings involving non-English languages.
@@ 2.2 Calibration of Post-training Pruning
Unlike sparse training (Yuan et al., 2021; Hoang and Liu, 2023; Zhang et al., 2023) or pruning-aware training (Liu et al., 2021), which require iterative training to achieve sparsity, post-training pruning (Frantar and Alistarh, 2023; Sun et al., 2024) does not require training but eliminates redundant weights based on the importance of weights that are calculated with calibration data. The retraining-free feature of post-training pruning makes it a more efficient approach for LLMs.
Prior research has primarily focused on calculating weight importance to optimize the performance of pruned models, examining importance metrics and importance ranking ranges (Frantar and Alistarh, 2023; Sun et al., 2024; Zhang et al., 2024). Two studies have investigated the impact of calibration data, looking at the quantity of calibration data (Zhang et al., 2024) and the sources of the data (Williams and Aletras, 2023). These studies have been limited to English. The effects of pruning on the multilingual capabilities of language models, and the impact of the language of calibration data on performance in other target languages remain unknown.
@@ 3 Methodology
To investigate the impact of the language of calibration data on the pruned model, we compare the performance (perplexity, and downstream tasks) in a range of languages between pruned models and their counterpart full-sized models, where different calibration data are applied.
3.1 Models
@@
Our experiments use two State-of-the-Art (SotA) LLM families: Llama-3 (Meta, 2024), the open-source SotA at the time of writing, and Aya-23 (Aryabumi et al., 2024), renowned for its multilingual pre-training. As our evaluation focuses on instructed generation tasks, also following Chrysostomou et al. (2023), we employ the instruction-tuned versions of both families. Consequently, our experimental setup includes Llama-3-instruction models in 8B and 70B parameter sizes, alongside Aya-instruction models in 8B and 35B parameter sizes.
3.2 Languages
@@
We include seven languages in our study: Arabic (AR), German (DE), English (EN), Spanish (ES), Russian (RU), Swahili (SW), and Chinese (ZH). This selection spans six language families, four writing systems, and encompasses both high and mid-resource languages. We include Swahili as an outlier calibration language, as neither Llama-3 or Aya-23 includes it in their pre-training corpora. A summary of languages used in our paper is given in Appendix A.
3.3 Pruning
Calibration
@@ 3.3 Pruning
We construct language-specific calibration sets from mC4 (Raffel et al., 2019) for the seven languages above. Specifically, following Frantar and Alistarh (2023); Sun et al. (2024), we randomly sample 128 sequences of 8192 tokens for each language, from the mC4 split for that language.111In the original implementation of Wanda and SparseGPT, each sample contains 2,048 tokens, matching the maximum context length of Llama 1. We increase the length to 8,192 tokens to match the maximum context of Llama-3. This is to avoid potential context length-dependence pruning behavior, although Sun et al. (2024) reports a neglectable return from increased calibration set sizes. For mixing different calibration languages, we mix in equal shares and keep the budget of 128 samples.
Pruning method
3.4 Evaluation Tasks
@@
We compare the performance of the pruned models calibrated on different languages using the perplexity on a subset of the mC4 validation set, and a selection of downstream tasks in different target languages: MMLU(Hendrycks et al., 2021), MKQA(Longpre et al., 2020), and Belebele(Bandarkar et al., 2023). To better isolate the impact of the calibration language, we also chose mARC and mHellaSwag(Dac Lai et al., 2023) to compare with their original versions ARC(Clark et al., 2018) and HellaSwag(Zellers et al., 2019) in English for their low sensitivity to the choice of calibration samples (Williams and Aletras, 2023). These tasks primarily assess commonsense reasoning, reading comprehension, and question answering using multiple choice questions. Unless stated otherwise, we evaluate in a zero-shot setting.
3.5 Implementation Details
@@
We adopt the code from Sun et al. (2024) for implementing model pruning. We use EleutherAI Evaluation Harness (Gao et al., 2024) for a robust and reproducible evaluation. We depend on Huggingface (Wolf et al., 2020) for loading datasets and models. All experiments are conducted with two NVIDIA A100 GPUs.
PPL | |||||||||
AR | DE | EN | ES | RU | SW | ZH | |||
Llama-3-8B-Instruct | - | - | 12.6 | 9.6 | 9.4 | 11.9 | 8.1 | 10.5 | 12.7 |
Wanda | AR | 26.3 | 17.3 | 15.9 | 20.0 | 14.0 | 30.4 | 24.9 | |
DE | 32.4 | 15.0 | 14.5 | 17.7 | 13.5 | 30.6 | 23.9 | ||
EN | 39.1 | 17.6 | 14.5 | 18.9 | 15.8 | 35.5 | 26.0 | ||
ES | 33.7 | 16.4 | 14.7 | 16.6 | 13.9 | 29.6 | 24.9 | ||
RU | 29.3 | 16.0 | 14.8 | 18.2 | 12.5 | 30.5 | 23.6 | ||
SW | 32.3 | 17.5 | 15.8 | 19.0 | 15.0 | 24.0 | 26.3 | ||
ZH | 37.9 | 20.1 | 17.2 | 21.2 | 16.8 | 38.9 | 24.4 | ||
SparseGPT | AR | 17.2 | 15.5 | 14.0 | 17.2 | 12.3 | 24.5 | 22.8 | |
DE | 30.7 | 12.1 | 13.3 | 15.8 | 11.9 | 25.5 | 23.6 | ||
EN | 34.4 | 14.7 | 13.0 | 17.3 | 13.6 | 28.6 | 23.1 | ||
ES | 32.4 | 14.3 | 13.6 | 14.2 | 13.0 | 26.3 | 25.7 | ||
RU | 27.4 | 14.5 | 13.8 | 16.7 | 9.9 | 25.6 | 22.8 | ||
SW | 24.8 | 15.7 | 13.9 | 17.0 | 14.2 | 14.0 | 25.5 | ||
ZH | 31.1 | 16.8 | 14.3 | 18.4 | 13.9 | 27.1 | 17.2 | ||
Aya-23-8B | - | - | 7.7 | 9.1 | 8.7 | 11.5 | 10.1 | 32.8 | 12.5 |
Wanda | AR | 11.1 | 13.6 | 11.5 | 16.2 | 15.4 | 53.8 | 19.5 | |
DE | 13.4 | 12.6 | 11.4 | 16.0 | 15.4 | 54.5 | 20.2 | ||
EN | 14.2 | 13.7 | 11.3 | 16.5 | 17.2 | 55.8 | 20.4 | ||
ES | 13.4 | 13.4 | 11.4 | 15.3 | 16.0 | 54.0 | 20.4 | ||
RU | 12.8 | 13.4 | 11.5 | 16.1 | 14.1 | 55.2 | 19.7 | ||
SW | 13.6 | 14.7 | 12.1 | 17.4 | 18.0 | 48.6 | 21.4 | ||
ZH | 13.4 | 14.4 | 11.8 | 17.3 | 17.1 | 57.2 | 18.4 | ||
SparseGPT | AR | 9.4 | 13.8 | 11.4 | 16.3 | 15.0 | 52.2 | 19.6 | |
DE | 13.4 | 11.4 | 11.1 | 15.3 | 14.4 | 53.8 | 20.3 | ||
EN | 14.6 | 13.3 | 10.8 | 16.2 | 17.4 | 54.7 | 20.7 | ||
ES | 13.2 | 12.7 | 11.1 | 13.9 | 16.1 | 53.3 | 20.6 | ||
RU | 12.7 | 13.1 | 11.3 | 16.0 | 12.2 | 54.6 | 19.8 | ||
SW | 11.7 | 13.7 | 11.5 | 16.2 | 16.3 | 36.9 | 21.1 | ||
ZH | 13.3 | 14.4 | 11.9 | 17.7 | 16.9 | 57.1 | 15.6 |

### v3

# Investigating Language-Specific Calibration For Pruning Multilingual Large Language Models

@@ Abstract
Recent advances in large language model (LLM) pruning have shown state-of-the-art (SotA) compression results in post-training and retraining-free settings while maintaining high predictive performance. However, previous research mainly considered calibrating based on English text, despite the multilingual nature of modern LLMs and their frequent use in non-English languages. In this paper, we set out to investigate calibrating the pruning of multilingual language models for monolingual applications. We present the first comprehensive empirical study, comparing different calibration languages for pruning multilingual models across diverse languages, tasks, models, and SotA pruning techniques. Our results offer practical suggestions, for example, calibrating in the target language can efficiently retain the language modeling capability but does not necessarily benefit downstream tasks. Through further analysis of latent subspaces, pruning masks, and individual neurons within pruned models, we find that while pruning generally preserves strong language-specific features, it may fail to retain language-specific neuron activation patterns and subtle, language-agnostic features associated with knowledge and reasoning that are needed for complex tasks.
@@ 1 Introduction
State-of-the-art language models often rely on over-parameterization with millions or billions of parameters, resulting in significant memory and computational demands (Zhang et al., 2017; Allen-Zhu et al., 2019; Xu & Du, 2023). Pruning is a model compression technique that removes unimportant weights to reduce memory footprint and inference computation (Gholami et al., 2022; Hoefler et al., 2021; Kuzmin et al., 2023). Recent pruning methods for language models, such as SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2024), demonstrated SotA performance in a post-training and retraining-free setting (Zhu et al., 2023). The pruning process involves passing a small number of examples, i.e. calibration data, to the model to determine the importance of weights for subsequent pruning (Kuzmin et al., 2022; Frantar & Alistarh, 2023; Kuzmin et al., 2023).
Notably, existing studies typically use calibration data in English and evaluate the post-pruning performance in English. On the other hand, most SotA LLMs are multilingual and frequently employed for tasks in non-English languages (Touvron et al., 2023; Achiam et al., 2023; Jiang et al., 2023). It remains unclear how to calibrate pruning to optimize the post-pruning performance of multilingual LLMs on tasks in non-English languages. In other words, the current literature provides little insight into efficient calibration strategies for multilingual LLMs targeting specific non-English languages. For example, if we aim to prune a multilingual LLM and use the pruned LLM for tasks in German, should we calibrate the compression process with text in English, German, or both?
This research, as the first, presents an extensive study investigating the impact of calibration data language on pruning LLMs for language-specific tasks. Further, we performed a set of analyses at the latent subspace level, the matrix level, and the neuron level, within the pruned models, to investigate the underlying causes for our observations. We summarize our key findings as follows:
Pruning generally affects language-agnostic features, which is potentially associated with reasoning and knowledge storage, more than it impacts language-specific features (Section 5.1); Pruning struggles to consistently identify essential neurons in the attention output and FFN down projection, potentially responsible for reasoning (Section 5.2); pruning also struggles to retain language-specific neurons with low activation probability (Section 5.3).
@@ 2.1 Multilingual Language Models
Most SotA LMs, such as Llama-3 (Meta AI, 2024) and Phi 3(Abdin et al., 2024), are trained on multilingual data, enabling them to understand and generate text in multiple languages (Huang et al., 2023; Holmström et al., 2023; Xu et al., 2024; Meta AI, 2024). Although these multilingual LMs follow the general training paradigm of training monolingual LMs, they are often found to behave differently or show unique features from monolingual models (Xu et al., 2024). For example, Deng et al. (2024) reveals that multilingual LMs are prone to generate unsafe outputs on malicious instructions in non-English languages, i.e. multilingual jailbreak. Wang et al. (2023) also identify that all LLMs tested in their study produce significantly more unsafe responses for non-English queries than English ones. Furthermore, previous work on model explanation finds that multilingual models are different from counterpart monolingual models in terms of their internal processes, i.e., different importance distributions on the same input (Jørgensen et al., 2022; Zhao & Aletras, 2024). A general conclusion drawn from these studies is that findings from monolingual settings, particularly those in English, are unlikely to hold in multilingual settings involving non-English languages.
@@ 2.2 Calibration of Post-training Pruning
Unlike sparse training (Yuan et al., 2021; Hoang & Liu, 2023; Zhang et al., 2023) or pruning-aware training (Liu et al., 2021), which require iterative training to achieve sparsity, post-training pruning (Frantar & Alistarh, 2023; Sun et al., 2024) does not require training but eliminates redundant weights based on the importance of weights that are calculated with calibration data. The retraining-free feature of post-training pruning makes it a more efficient approach for LLMs.
Prior research has primarily focused on calculating the importance of weight to optimize the performance of pruned models, examining importance metrics and importance ranking ranges (Frantar & Alistarh, 2023; Sun et al., 2024; Zhang et al., 2024). Two studies have investigated the impact of calibration data, looking at the quantity of calibration data (Zhang et al., 2024) and the sources of the data (Williams & Aletras, 2023). These studies have been limited to English. The effects of pruning on the multilingual capabilities of language models, and the impact of the language of calibration data on performance in other target languages remain unknown.
@@ 3 Methodology
To investigate the impact of the language of calibration data on the pruned model, we compare the performance (pruning error, Signal-to-noise ratio, perplexity, and downstream tasks) in a range of languages between pruned models and their counterpart full-sized models, where different calibration data are applied.
@@
Models Our experiments use two SotA LLM families: Llama-3 (Meta, 2024), the open-source SotA at the time of writing, and Aya-23 (Aryabumi et al., 2024), renowned for its multilingual pre-training.As our evaluation focuses on instructed generation tasks, we employ the instruction-tuned versions of both families (Chrysostomou et al., 2023). Consequently, our experimental setup includes Llama-3-instruction models in 8B and 70B parameter sizes, alongside Aya-instruction models in 8B and 35B parameter sizes.
@@
Languages We include seven languages in our study: Arabic (AR), German (DE), English (EN), Spanish (ES), Russian (RU), Swahili (SW), and Chinese (ZH). This selection spans six language families, four writing systems, and encompasses both high and mid-resource languages. We include Swahili as an outlier calibration language, as neither Llama-3 or Aya-23 includes it in their pre-training corpora. A summary of languages used in our paper is given in Table 5 in Appendix B.
@@
Pruning and Calibration We construct calibration sets in seven different languages from mC4 (Raffel et al., 2019). Specifically, following Frantar & Alistarh (2023); Sun et al. (2024), we randomly sample 128 sequences of 8192 tokens for each language, from the mC4 split for that language.111In the original implementation of Wanda and SparseGPT, each sample contains 2,048 tokens, matching the maximum context length of Llama 1. We increase the length to 8,192 tokens to match the maximum context of Llama-3. This is to avoid potential context length-dependence pruning behavior, although Sun et al. (2024) reports a negligible return from increased calibration set sizes. For mixing different calibration languages, we mix in equal shares and keep the budget of 128 samples.
We focus on two SotA post-training pruning methods, Wanda (Sun et al., 2024) and SparseGPT (Frantar & Alistarh, 2023). Unless stated otherwise, we prune for 50% unstructured sparsity to impose fewer restrictions while keeping all other hyperparameters as in the original paper.
@@
Evaluation Downstream Tasks We compare the performance of the pruned models calibrated on different languages using the perplexity on a subset of the mC4 validation set, and a selection of downstream tasks in different target languages: MMLU (Hendrycks et al., 2021), MKQA (Longpre et al., 2020), and Belebele (Bandarkar et al., 2023). To better isolate the impact of the calibration language, we also chose mARC and mHellaSwag (Dac Lai et al., 2023) to compare with their original versions, ARC (Clark et al., 2018) and HellaSwag (Zellers et al., 2019) in English for their low sensitivity to the choice of calibration samples (Williams & Aletras, 2023). These tasks primarily assess commonsense reasoning, reading comprehension, and question answering using multiple choice questions. We present further details of each task in Appendix C. Unless stated otherwise, we evaluate in a zero-shot setting.
PPL | SNR | Pruning Error | |||||||||||||||||||||
AR | DE | EN | ES | RU | SW | ZH | AR | DE | EN | ES | RU | SW | ZH | AR | DE | EN | ES | RU | SW | ZH | |||
Llama-3-8B-Instruct | Wanda | AR | 26.3 | 17.3 | 15.9 | 20.0 | 14.0 | 30.4 | 24.9 | 7.0 | 7.8 | 7.2 | 7.1 | 8.5 | 5.9 | 9.7 | 85.0 | 87.0 | 100.6 | 86.2 | 119.6 | 113.9 | 86.7 |
DE | 32.4 | 15.0 | 14.5 | 17.7 | 13.5 | 30.6 | 23.9 | 6.5 | 8.4 | 8.0 | 7.3 | 8.5 | 5.8 | 9.7 | 94.1 | 77.0 | 77.7 | 81.4 | 130.1 | 117.0 | 89.3 | ||
EN | 39.1 | 17.6 | 14.5 | 18.9 | 15.8 | 35.5 | 26.0 | 6.1 | 7.6 | 8.2 | 7.1 | 8.3 | 5.3 | 10.6 | 103.5 | 92.8 | 76.0 | 85.0 | 127.1 | 130.6 | 62.6 | ||
ES | 33.7 | 16.4 | 14.7 | 16.6 | 13.9 | 29.6 | 24.9 | 6.3 | 7.7 | 8.0 | 7.4 | 8.2 | 5.7 | 9.2 | 98.8 | 92.0 | 79.3 | 81.8 | 139.0 | 119.3 | 107.7 | ||
RU | 29.3 | 16.0 | 14.8 | 18.2 | 12.5 | 30.5 | 23.6 | 6.8 | 8.1 | 8.1 | 7.5 | 10.4 | 5.8 | 10.0 | 87.3 | 82.6 | 76.5 | 77.8 | 66.2 | 119.1 | 84.3 | ||
SW | 32.3 | 17.5 | 15.8 | 19.0 | 15.0 | 24.0 | 26.3 | 6.5 | 7.7 | 7.5 | 7.0 | 8.3 | 6.7 | 9.6 | 92.8 | 89.9 | 88.3 | 87.2 | 127.3 | 94.9 | 87.0 | ||
ZH | 37.9 | 20.1 | 17.2 | 21.2 | 16.8 | 38.9 | 24.4 | 6.0 | 7.1 | 7.7 | 6.6 | 8.0 | 4.9 | 10.0 | 105.4 | 103.2 | 83.1 | 94.0 | 131.2 | 142.5 | 87.3 | ||
SparseGPT | AR | 17.2 | 15.5 | 14.0 | 17.2 | 12.3 | 24.5 | 22.8 | 11.0 | 9.0 | 8.7 | 8.6 | 11.2 | 7.3 | 11.8 | 34.7 | 65.4 | 67.8 | 58.9 | 52.9 | 84.1 | 46.4 | |
DE | 30.7 | 12.1 | 13.3 | 15.8 | 11.9 | 25.5 | 23.6 | 7.4 | 11.6 | 8.1 | 9.2 | 11.6 | 7.2 | 11.8 | 76.0 | 37.3 | 118.6 | 52.6 | 48.8 | 87.7 | 46.1 | ||
EN | 34.4 | 14.7 | 13.0 | 17.3 | 13.6 | 28.6 | 23.1 | 7.0 | 9.4 | 11.4 | 8.6 | 11.0 | 6.6 | 11.9 | 83.6 | 59.9 | 37.4 | 59.9 | 57.0 | 99.3 | 46.0 | ||
ES | 32.4 | 14.3 | 13.6 | 14.2 | 13.0 | 26.3 | 25.7 | 7.4 | 9.6 | 8.5 | 10.9 | 11.4 | 7.1 | 11.6 | 76.0 | 57.6 | 83.6 | 36.4 | 52.7 | 89.8 | 48.7 | ||
RU | 27.4 | 14.5 | 13.8 | 16.7 | 9.9 | 25.6 | 22.8 | 7.8 | 9.5 | 8.7 | 8.9 | 13.8 | 7.1 | 12.1 | 70.4 | 58.3 | 73.6 | 55.6 | 31.1 | 89.0 | 43.0 | ||
SW | 24.8 | 15.7 | 13.9 | 17.0 | 14.2 | 14.0 | 25.5 | 8.0 | 9.0 | 8.2 | 8.5 | 10.8 | 11.4 | 11.1 | 67.5 | 66.3 | 80.9 | 61.7 | 59.5 | 33.7 | 54.7 | ||
ZH | 31.1 | 16.8 | 14.3 | 18.4 | 13.9 | 27.1 | 17.2 | 7.2 | 8.6 | 8.6 | 8.0 | 10.8 | 6.6 | 15.7 | 79.9 | 71.5 | 68.3 | 68.4 | 60.3 | 97.9 | 19.5 | ||
Aya-23-8B | Wanda | AR | 11.1 | 13.6 | 11.5 | 16.2 | 15.4 | 53.8 | 19.5 | 6.2 | 6.4 | 6.8 | 6.4 | 6.5 | 6.1 | 5.8 | 676.7 | 1033.2 | 851.7 | 929.5 | 911.4 | 1023.1 | 420.0 |
DE | 13.4 | 12.6 | 11.4 | 16.0 | 15.4 | 54.5 | 20.2 | 6.0 | 6.7 | 6.9 | 6.6 | 6.6 | 6.2 | 5.8 | 656.4 | 992.2 | 829.0 | 897.5 | 888.2 | 994.9 | 418.2 | ||
EN | 14.2 | 13.7 | 11.3 | 16.5 | 17.2 | 55.8 | 20.4 | 5.9 | 6.5 | 7.0 | 6.5 | 6.4 | 6.1 | 5.7 | 660.8 | 997.2 | 829.8 | 903.1 | 903.3 | 1001.8 | 426.1 | ||
ES | 13.4 | 13.4 | 11.4 | 15.3 | 16.0 | 54.0 | 20.4 | 6.1 | 6.6 | 7.0 | 6.8 | 6.6 | 6.3 | 5.8 | 646.2 | 979.3 | 819.5 | 873.8 | 869.0 | 973.3 | 411.4 | ||
RU | 12.8 | 13.4 | 11.5 | 16.1 | 14.1 | 55.2 | 19.7 | 6.2 | 6.7 | 7.0 | 6.8 | 6.8 | 6.4 | 5.9 | 636.5 | 955.9 | 808.8 | 853.0 | 859.1 | 951.0 | 404.5 | ||
SW | 13.6 | 14.7 | 12.1 | 17.4 | 18.0 | 48.6 | 21.4 | 5.9 | 6.3 | 6.7 | 6.3 | 6.3 | 6.3 | 5.6 | 680.2 | 1049.4 | 850.9 | 949.4 | 932.0 | 1043.1 | 430.3 | ||
ZH | 13.4 | 14.4 | 11.8 | 17.3 | 17.1 | 57.2 | 18.4 | 5.9 | 6.3 | 6.8 | 6.4 | 6.3 | 6.0 | 6.1 | 684.8 | 1036.2 | 842.0 | 923.0 | 930.2 | 1030.3 | 416.3 | ||
SparseGPT | AR | 9.4 | 13.8 | 11.4 | 16.3 | 15.0 | 52.2 | 19.6 | 17.1 | 15.1 | 14.7 | 12.8 | 15.6 | 15.3 | 11.2 | 30.7 | 69.6 | 75.5 | 135.4 | 55.0 | 64.1 | 79.0 | |
DE | 13.4 | 11.4 | 11.1 | 15.3 | 14.4 | 53.8 | 20.3 | 11.7 | 17.9 | 15.6 | 14.4 | 16.1 | 15.5 | 9.0 | 108.2 | 36.8 | 59.4 | 83.9 | 49.5 | 62.2 | 153.3 | ||
EN | 14.6 | 13.3 | 10.8 | 16.2 | 17.4 | 54.7 | 20.7 | 11.8 | 16.2 | 17.9 | 13.9 | 15.4 | 15.4 | 9.9 | 101.7 | 54.0 | 33.7 | 93.0 | 57.4 | 62.8 | 114.6 | ||
ES | 13.2 | 12.7 | 11.1 | 13.9 | 16.1 | 53.3 | 20.6 | 12.0 | 16.1 | 15.3 | 18.0 | 15.8 | 15.6 | 9.1 | 101.4 | 55.7 | 65.1 | 34.2 | 52.1 | 59.6 | 147.2 | ||
RU | 12.7 | 13.1 | 11.3 | 16.0 | 12.2 | 54.6 | 19.8 | 12.2 | 15.3 | 14.2 | 13.6 | 18.4 | 15.2 | 9.4 | 97.8 | 68.1 | 91.3 | 104.2 | 30.1 | 66.8 | 139.9 | ||
SW | 11.7 | 13.7 | 11.5 | 16.2 | 16.3 | 36.9 | 21.1 | 12.4 | 15.7 | 15.2 | 14.0 | 15.4 | 19.5 | 8.5 | 93.4 | 60.2 | 63.4 | 91.9 | 58.1 | 26.3 | 178.2 | ||
ZH | 13.3 | 14.4 | 11.9 | 17.7 | 16.9 | 57.1 | 15.6 | 12.5 | 14.7 | 13.9 | 13.3 | 15.1 | 14.7 | 16.8 | 86.2 | 78.7 | 95.3 | 107.4 | 61.4 | 76.1 | 22.4 |
@@
Implementation Details We adopt the code from Sun et al. (2024) for implementing model pruning. We use EleutherAI Evaluation Harness (Gao et al., 2024) for a robust and reproducible evaluation. We use Huggingface (Wolf et al., 2020) for loading datasets and models. All experiments are conducted with at most two NVIDIA A100 GPUs.
4 Results
4.1 Pruning Results
Table 1 presents the perplexity (PPL), Signal-to-noise ratio (SNR) and pruning error of hidden states for models pruned to 50% sparsity and calibrated on different languages. SNR and pruning error estimate the pruning accuracy (pruning performance); ideal pruning is able to preserve activations identical to the full model. The formal expression and detailed explanation of SNR and pruning error are presented in the Appendix D.1. Perplexity reflects the general language modeling capability. The column headers denote the evaluation languages, while the row headers indicate the calibration languages.
Overall, our findings reveal that no single calibration language consistently outperforms others across perplexity, pruning error, or SNR metrics. The optimal calibration language varies depending on the target evaluation language. Notably, calibrating on the target language itself generally yields the best pruning performance and the lowest perplexity, as evidenced by the diagonal pattern in Table 1. There are a few exceptions. For instance, when evaluating Llama-3 pruned with Wanda on Chinese, Russian calibration performs optimally on perplexity (23.6), slightly outperforming Chinese calibration (24.4).
4.2 Downstream Task Performance